6. WHAT IS THE TEST STATISTIC?

A manufacturer claims that 10% of women using the "pill" suffer from side effects. The Federal Drug Administration (FDA) believes that the manufacturer's claim is too low and decides to test the manufacturer's claim at α = 5 %. A random sample of 900 women who use the manufacturer's pill shows that 103 suffer side effects.
Answer #1

Given that a manufacturer claims that 10% (p = 0.10) of women using the "pill" suffer from side effects. The Federal Drug Administration (FDA) believes that the manufacturer's claim is too low and decides to test the manufacturer's claim at α = 5 %. A random sample of n = 900 women who use the manufacturer's pill shows that X =103 suffer side effects.

Now based on the claim the hypotheses are:

Ho: p = 0.10

Ha: p > 0.10

the sample proportion is calculated as:

\bar{p} = \frac{X}{n} = \frac{103}{900} = 0.1144

Test statistic:

The test statistic is calculated as:

2 = = P - po po(1 - pon 0.1144 - 0.10 0.10(1 - 0.10/900 = 1.444